OpenStudy (anonymous):

solve each equation system by the addition method 6x-10y=2 2x+5y=9

OpenStudy (anonymous):

Before you add both systems together you should multiply the second equation by -3 to eliminate x. When you do that you should have : 6x -10y = 2 -6x - 15y = -27

OpenStudy (anonymous):

Now when you add this is what you get: -25y = -25 Therefore y = 1

OpenStudy (anonymous):

If y = 1 we will substitute this value into either original equation 1 or 2. So lets pick equation 1: 6x - 10y = 2 if y = 1 we have 6x - 10 = 2 6x = 12 x = 2 So your final solution in ordered pair: (2,1)